OpenStudy (anonymous):

Solve for d. 4d + 7 = -17 A. d = -6 B. d = -2.5 C. d = 2.5 D. d = -4.25

OpenStudy (anonymous):

4d =-17-7 4d=-24 d=-24/4 d=-3

OpenStudy (mathstudent55):

First, subtract 7 from both sides. Then divide both sides by 4.

OpenStudy (mathstudent55):

@heybhai Your algebra is good, but you need to check your arithmetic.

OpenStudy (anonymous):

sorry its -6 not -3

OpenStudy (uri):

Subract 7 both sides. 4d + 7 = -17 -7 -7 ========== 4d =-24

OpenStudy (anonymous):

@uri -24/4 is -6